Analysts forecast that a company's dividends in 2019, 2020 and 2021 will be the same amount as the dividends paid in 2018 which was 0.05. After 2021 dividends will increase by 2% p.a. What is a fair price for a share in 2019? Required rate of return is 9%
Answer #1

Price of a stock is the present value of all future cash flows receivable from the stock discounted at required rate of return

Price of the stock at the end of 2021

= Dividend of 2021 x (1 + Growth) / (Required rate of return – Growth rate)

= $0.05 x 1.02 / (0.09 – 0.02)

= $0.7286

So, price of the stock today is calculated in the following table

Present value factor

= 1 / (1 + r) ^ n

Where,

r = Rate of return = 9% or 0.09

n = Years 1 to 3

So, PV Factor for year 2 will be

= 1 / (1.09^2)

= 1 / 1.1881

= 0.841680

Calculations A B C = A x B
Years Cash Flow PV Factor Present Value
1 0.05 0.917431 0.04587156
2 0.05 0.84168 0.042084
3 0.05 0.772183 0.038609174
3 0.7286 0.772183 0.562612884
Price 0.6892

So, price today is $0.6892